OpenStudy (anonymous):

What is the slope of the line that is perpendicular to the line whose equation is 2x + y = 4.

OpenStudy (vhtran):

y=mx+b Try to put it in that format

OpenStudy (vhtran):

y=-2x+4 I moved it so that y is on the left and x and the constant are on the right

OpenStudy (vhtran):

In y=mx+b "m" stands for the slope so the slope for that line is -2

OpenStudy (vhtran):

the slope of the line that is perpendicular is just the reciprocal of -2.

OpenStudy (anonymous):

The gradient of a line is the value of m when you look at the equation y=mx+c, so for this example it is 1/2. The perpendicular gradient is given as -1/m, so you can work out the perpendicular gradient is -2.